OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (anonymous):

will i need identities for this

OpenStudy (anonymous):

This one is somewhat trickier. You have to use the Pythagorean identity: \[\cos^2\theta+\sin^2\theta=1\] Notice that \(x^2=9\cos^2t\) and \(y^2=9\sin^2t\), and that adding them together you have \[x^2+y^2=9\left(\cos^2t+\sin^2t\right)\] Utilizing the identity, you get \[x^2+y^2=9\]